OpenStudy (anonymous):

Find the quotient. 6/2/7÷ 2/3/4

OpenStudy (anonymous):

I ended up getting 0.01785714285 on calculator that's not the way I want it it should be like above

OpenStudy (solomonzelman):

Ok, here, use this to put each mixed fraction, as an improper fraction.\[\huge\color{blue}{ a \frac{b}{c} =\frac{(a \times c)+b}{c} } \] after you have them in a d/f form, then you divide them, lets say you are dividing\[\huge\color{blue}{ \frac{h}{j} \div\frac{r}{t} } \] you flip the fraction you are dividing by and multiply by it,\[\frac{h}{j} \times\frac{t}{r}\]

OpenStudy (texaschic101):

turn your mixed numbers into improper fractions. Can you do that ?

OpenStudy (mathstudent55):

Can you draw this problem? I don't understand what you mean by 6/2/7

OpenStudy (texaschic101):

she means 6 2/7

OpenStudy (texaschic101):

I am not sure why she writes it like she does

OpenStudy (mathstudent55):

This is the problem?